OpenStudy (anonymous):

Make G the subject of the formula. 3e(g+4)/ 7-g =2

OpenStudy (anonymous):

3E(G+4)/7-G =2 3E(G+4)=14-2G 3EG+12E=14-2G (3E+2) G=14-12E G= (14-12E)/(3E+2)

Parth (parthkohli):

In other words, we have to isolate g. 1) Multiply 7 - g to both sides. 2) Take g on one side and constants on the other. 3) Solve for g.

Parth (parthkohli):

Got it?

OpenStudy (diyadiya):

You're Right! @Courtney-Leigh

OpenStudy (anonymous):

Thank you:D